# Semi-analytical covariance matrices for two-point correlation function for DESI 2024 data

### v2

@@ Abstract
We present an optimized way of producing the fast semi-analytical covariance matrices for the Legendre moments of the two-point correlation function, taking into account survey geometry and mimicking the non-Gaussian effects. We validate the approach on simulated (mock) catalogs for different galaxy types, representative of the Dark Energy Spectroscopic Instrument (DESI) Data Release 1, used in 2024 analyses. We find only a few percent differences between the mock sample covariance matrix and our results, which can be expected given the approximate nature of the mocks, although we do identify discrepancies between the shot-noise properties of the DESI fiber assignment algorithm and the faster approximation used in the mocks. Importantly, we find a close agreement ( relative differences) in the projected errorbars for distance scale parameters for the baryon acoustic oscillation measurements. This confirms our method as an attractive alternative to simulation-based covariance matrices, especially for non-standard models or galaxy sample selections, in particular, relevant to the broad current and future analyses of DESI data.
@@ 1 Introduction
It is a particularly exciting time for observational cosmology due to the transition from Stage III to Stage IV dark energy experiments. The Dark Energy Spectroscopic Instrument (DESI) [1, 2] belongs to this newer generation and is actively operating. Last year saw the validation of its scientific program [3] and the early data release [4]. As we are writing this paper in 2024, a larger 1-year dataset (DR1) [5] is being released, along with two-point clustering [6], inverse distance ladder measurements (and thus the expansion history of the Universe) using the baryon acoustic oscillations (BAO) of galaxies, quasars [7], and Lyman- [8], full-shape analysis of the 2-point statistics for galaxies and quasars constraining the growth of cosmic structure [9], and implications for cosmological models [10, 11, 12].
@@
For meaningful interpretation of the data, covariance matrices are necessary. The gold standard in large-scale structure studies has been based on suites of simulated (mock) catalogs (e.g. [13, 14, 15, 16, 17]). They need to be both highly accurate representations of the data (in particular, large enough to cover the survey volume), and numerous enough to give a good estimate of the covariance matrix of the vector of observables (since the relative precision is primarily determined by the number of samples and the dimension of the matrix). Since highly detailed simulations require a lot of time even for smaller volumes, it is unavoidable to rely on approximations, limiting the realism of the simulated catalogs. Even with that, the generation and calibration of an adequate mock suite are very hard and expensive.
@@
With time, this approach is becoming only more challenging. First, as the surveys improve, each mock catalog needs to include more galaxies and/or more volume, thereby taking longer to generate and process. Second, with more data, we aim to include a longer vector of observables in the analysis, requiring a larger covariance matrix for it, which in turn demands a higher number of mocks for adequate precision [18, 19, 20]. Third, the substantial time to generate mocks typically means that one cannot produce enough simulations for many separate sets of cosmological, galaxy-halo connection models and selections of tracer galaxy samples. This creates a potential systematic error when extrapolating the covariances derived in one scenario to another. Fourth, as a specific example of this, the long timeframe to generate mocks can even create a situation where schedule concerns force the mocks to be calibrated on early inputs that do not match the final version of the analysis. The problem is especially severe when blinding is employed, and the simulation teams should not see the true and complete data clustering before the analysis methodology is frozen.
This creates a need and opportunity for faster alternative methods. First, we note the development of analytical covariance matrices for galaxy power spectrum using perturbation theory [21, 22, 23]. Unfortunately, perturbative expansions of the higher-point functions do not translate well to the correlation functions due to the nonlocality of the Fourier transform connecting them with power spectra.
@@
Second, we highlight the jackknife technique, which involves the re-sampling of the data to estimate the covariance. In the context of galaxy surveys, it is challenging to divide the volume into equivalent parts because of anisotropic effects including those of boundaries, the harder the more regions one attempts to create. The number of samples still impacts the precision of the covariance matrix estimate, like with the mocks. Moreover, the parts will not be independent from each other. [24] attempted to correct for the last factor by dividing the jackknife pair counts into different categories and re-weighting some of them. However, [25] show that some of the former assumptions are violated in higher-density setups leading to significant biases. To solve this problem and obtain a more precise, better-conditioned matrix, they propose a hybrid approach, combining jackknife with mocks (requiring not as many of them as for the sample covariance).
@@
For this work, we choose an approach developed in a series of papers [26, 27, 28, 29] and implemented as the RascalC code111https://github.com/oliverphilcox/RascalC. It focuses on configuration space (correlation functions) and enables the computation of covariance matrices using only the data, without any mocks, by employing both analytical and jackknife methods. These works are reviewed in more detail in Section 2.1.
@@
The abovementioned approach has been successfully applied and validated for the covariance matrices of the early DESI data [30, 31]. Following that, it was embraced as part of a coordinated covariance matrix effort for DESI DR1 two-point clustering measurements, along with analytical covariance matrices for power spectra [23] and the general comparison focusing on the consistency of the model fits [32]. We have benefited enormously from synergies with other supporting studies for the galaxies and quasars BAO [7]: optimal reconstruction [33, 34], combined tracers [35], halo occupation distribution systematics [36, 37], fiducial cosmology systematics [38] and theoretical systematics [39].
@@ 2.1 Estimation of covariance matrix for the two-point correlation functions
We start by briefly reviewing the methodology developed in [26, 27, 28, 31]. The RascalC code builds single-parameter222For single tracer; for multiple tracers it is one parameter per tracer. covariance matrix models333This step is the most computationally heavy and is implemented in C++. based on a random catalog and a table of the 2-point correlation function values. Then we fit a model to the reference covariance to obtain the optimal parameter value for the final prediction. In the fiducial data pipeline, shown schematically in Fig. 1, we measure the correlation function directly from the data, the code produces separate models for full and jackknife covariance matrices, we fit the latter to the data jackknife covariance matrix, and plug the resulting optimal parameter into the full model. In an alternative pipeline (Appendices C and 4), we use the best fit of the full covariance model to the mock sample covariance instead. In the following, we provide more details.
@@
The Landy-Szalay estimator [40] for the 2-point correlation function (2PCF) in radial bin and angular bin444RascalC assumes uniform binning in , being the cosine of the angle between the line of sight and the pair separation (assuming symmetry with respect to ). The line of sight is midpoint in aperiodic survey and in periodic boxes. is
where , are random points and are data points; one can further write
@@
Here a grid of cubic cells has been assigned to the survey such that each cell contains no more than one galaxy, is the expected mean number density of tracer in the cell , is the expected mean weight, is the fractional galaxy overdensity, is the absolute value of the cosine of the angle between the line of sight and the separation vector ( being its absolute value), and are binning functions (unity if the argument fits into the bin and zero otherwise).
Then the covariance matrix can be computed, for convenience designated as
The covariance can be expanded using Eqs. 2.1 and 2.2 and separated into sums over configurations of 4, 3, and 2 points (since some of the four can coincide) [26]. The shot-noise approximation was then used to deal with squares of overdensity:
It includes a shot-noise rescaling parameter . The “default”, Poissonian shot-noise corresponds to , but the parameter can be set to other values for different tracers to improve the realism of the covariance matrix, as detailed below. An intuitive expectation is for , since non-Gaussianity typically enhances the small-scale clustering, although it has not been proven strictly.
Previous works have established the procedure of taking the Gaussian limit: nulling the 3-point function and the connected 4-point function, and instead adjusting the covariance matrix with shot-noise rescaling parameter(s) according to Eq. A.1. This does not remove the 3- and 4-point terms completely, because they have contributions from the 2PCF.
@@
The 4, 3, and 2-point terms (Eq. A.2) are estimated using Monte Carlo importance sampling of points from random catalogs [28]; the RascalC code also requires random counts for normalization and a table of correlation function values to evaluate the summands in Eq. A.2 [31].
The shot-noise rescaling parameter can be optimized to match a sample covariance based on a smaller set of mocks [26] or to jackknife covariance from the data itself [27]. The latter only requires mocks for initial validation (like in this paper), and then allows to generate covariance matrices for different setups more quickly and easily, without the need to generate or adjust simulations for them. We advise the reader to remark that the code does not fit the final covariance to the jackknife one, but makes a separate theoretical prediction for the latter, taking into account the correlations between the jackknife regions.
@@
The code uses a slightly non-standard formalism, dubbed unrestricted jackknife [28]. There the jackknife correlation function estimate is not the auto-correlation of the whole survey excluding the jackknife region , but the cross-correlation function between that region and the whole survey. Equivalently, this means that the additional jackknife weighting factor for the pair of points , , is 1 if both of them belong to the jackknife region , if only one and 0 if neither. Then the pair counts can be converted from different terms (auto and cross jackknife counts) often saved separately in light of Mohammad-Percival correction [24].
@@
The unrestricted jackknife is convenient since the full pair counts of all types are the sum of all the jackknife ones. So if one weighs the regions by the pair counts,
@@
the weighted mean correlation function is equal to the full-survey one.
@@
the corresponding theoretical estimate, , is constructed analogously to Eq. A.1 but with different terms defined in Eq. A.3.
Shot-noise rescaling has been obtained for each tracer separately by fitting the prediction for its auto-correlation function’s jackknife covariance (constructed analogously to Eq. 2.5 but with the terms given by Eq. A.3) to the data (Eq. 2.7). More specifically, this involves the minimization of the Kullback-Leibler (KL) divergence, where the RascalC covariance is inverted [28] (since the jackknife one is often not invertible). The final covariance is obtained by plugging the resulting shot-noise rescaling values into Eq. A.1.
Theoretically, inversion of the RascalC covariance gives a slightly biased estimate of the precision matrix. However, the Hartlap factor [18] is not applicable since it is not a sample covariance. The relevant correction matrix accounting for importance sampling noise has been worked out [27], but we find it practically insignificant: the eigenvalues deviate from 1 by .
Standard BAO reconstruction procedures shift the positions of both the data and random points. Original random points are kept as well, the shifted ones are commonly denoted by . The correlation function is estimated via the Landy-Szalay estimator (Eq. 2.1) with instead of . As [31] argued, in the computations of covariance matrices for reconstructed catalogs we used the shifted randoms in the importance sampling, not shifted random counts for normalization in Eqs. A.2 and A.3 and slightly differently normalized two-point correlation function for the integrands:
To summarise, the main approximation is that shot-noise rescaling of purely Gaussian contributions (i.e., ignoring 3-point and connected 4-point function, only based on 2-point) can produce a realistic covariance matrix in configuration space. A theoretical motivation for this is that non-Gaussian contributions primarily affect the squeezed configurations involving small-scale correlations, below the bin width for the 2-point function, while the shot noise regulates infinitesimally small scales. The method has been empirically shown to agree well enough with the mock-based covariances [26, 41, 31] so far.

### v3

@@ Abstract
We present an optimized way of producing the fast semi-analytical covariance matrices for the Legendre moments of the two-point correlation function, taking into account survey geometry and mimicking the non-Gaussian effects. We validate the approach on simulated (mock) catalogs for different galaxy types, representative of the Dark Energy Spectroscopic Instrument (DESI) Data Release 1, used in 2024 analyses. We find only a few percent differences between the mock sample covariance matrix and our results, which can be expected given the approximate nature of the mocks, although we do identify discrepancies between the shot-noise properties of the DESI fiber assignment algorithm and the faster approximation (emulator) used in the mocks. Importantly, we find a close agreement ( relative differences) in the projected errorbars for distance scale parameters for the baryon acoustic oscillation measurements. This confirms our method as an attractive alternative to simulation-based covariance matrices, especially for non-standard models or galaxy sample selections, making it particularly relevant to the broad current and future analyses of DESI data.
@@ 1 Introduction
It is a particularly exciting time for observational cosmology due to the transition from Stage III to Stage IV dark energy experiments. The Dark Energy Spectroscopic Instrument (DESI) [1, 2] belongs to this newer generation and is actively operating. Last year saw the validation of its scientific program [3] and the early data release [4]. As we are writing this paper in 2024, a larger 1-year dataset (DR1) [5] is being released, along with two-point clustering [6], inverse distance ladder measurements (and thus the expansion history of the Universe) using the baryon acoustic oscillations (BAO) of galaxies, quasars [7], and Lyman- [8], full-shape analysis of the 2-point statistics for galaxies and quasars constraining the growth of cosmic structure [9], and implications for cosmological models [10, 11, 12]. These DESI results have unprecedented precision for their kind of measurement, providing unique new opportunities to test our understanding of the Universe.
Data interpretation with a physical model requires a covariance matrix model, which can be hard to obtain. An intuitive way to do so is from the scatter in repeated independent identical measurements. However, only one Universe is accessible for us to probe in cosmology. Moreover, it is not feasible to replicate a state-of-the-art experiment exactly. Therefore covariance matrix estimation in cosmology requires elaborate techniques.
@@
The standard method for computing covariance matrices in large-scale structure studies has been based on scatter within large sets of simulated (mock) catalogs (e.g. [13, 14, 15, 16, 17]). They need to be both highly accurate representations of the data (in particular, large enough to cover the survey volume), and numerous enough to give a good estimate of the covariance matrix of the vector of observables. The relative precision is primarily determined by the number of samples and the dimension of the matrix. If the number of samples is smaller than the number of bins plus one, the resulting covariance matrix estimate is not invertible. Moreover, sample noise biases the estimate of the inverse covariance matrix [18] and causes the widening of model parameter errorbars [19]. Because highly detailed simulations require a lot of time even for volumes much smaller than a survey like DESI covers, it is unavoidable to rely on approximations, limiting the realism of the simulated catalogs. Even so, generating and calibrating an adequate mock suite is very hard and expensive.
@@
The mock-based covariance matrix estimation is becoming only more challenging with time. First, as the surveys improve, each mock catalog needs to include more galaxies and/or more volume, thereby taking longer to generate and process. Second, with more data, we aim to include a longer vector of observables in the analysis, requiring a larger covariance matrix for it, which in turn demands a higher number of mocks for adequate precision [18, 20, 21]. Third, the substantial time to generate mocks typically means that one cannot produce enough simulations for many separate sets of cosmological, galaxy-halo connection models and selections of tracer galaxy samples. This creates a potential systematic error when extrapolating the covariances derived in one scenario to another. Fourth, as a specific example of this, the long timeframe to generate mocks can even create a situation where schedule concerns force the mocks to be calibrated on early inputs that do not match the final version of the analysis. The problem is especially severe when blinding is employed, and the simulation teams should not see the true and complete data clustering before the analysis methodology is frozen. This creates a need and opportunity for faster alternative methods.
A promising alternative is a theoretical derivation of the covariance matrix, which is more conveniently performed in Fourier space for cosmological clustering. Analytical covariance matrices for galaxy power spectrum have been developed using perturbation theory [22, 23, 24]. However, perturbative expansions only work for a limited range of wavenumbers. Unfortunately, the translation of these results to the configuration space introduces an additional precision loss due to the nonlocality of the Fourier transform. The accuracy of the transformation can be improved with the generalization of the FFTLog algorithm to covariance matrices [25], but this approach has not yet been applied to and validated for redshift-space clustering of galaxies in 3 dimensions. Therefore the analytical methods are not yet directly applicable to the correlation functions.
@@
Another approach is re-sampling the data and considering scatter between different parts for the covariance estimation. The jackknife technique is based on this idea. Unfortunately, dividing the spectroscopic survey volume into equivalent parts is challenging because of boundary effects and multiple other factors varying across the survey. This problem becomes more severe with a larger number and accordingly smaller size of regions. More samples are still desirable for higher covariance matrix precision, analogously to the mock case. In addition, regions are not completely independent from each other. [26] attempted to correct for the last factor by dividing the jackknife pair counts into different categories and re-weighting some of them. However, [27] show that some of the former assumptions are violated in higher-density setups leading to significant biases. To solve this problem and obtain a more precise, better-conditioned matrix, they propose a hybrid approach, combining jackknife with mocks (requiring fewer realizations than for the sample covariance). This illustrates the promise of hybrid approaches incorporating different techniques.
@@
Here we focus on a combination of analytical and jackknife elements in configuration space. The theoretical component hinges on the relation between the covariance matrix of the 2-point correlation function (2PCF) and the 4-, 3- and 2-point functions. This method has been developed in a series of papers [28, 29, 30, 31] and implemented as the RascalC code111https://github.com/oliverphilcox/RascalC. This approach enables the computation of covariance matrices for correlation functions using only the data, without extra effort, assumptions and approximations for generating mocks.
Multipole (Legendre) moments of the correlation function are preferable for redshift-space analysis and theoretical modeling (e.g. [32]). [30] only developed an estimator for the covariance matrix of the 2PCF in angular bins. With a large number of angular bins required to estimate multipoles adequately, the computation time of the covariance matrix becomes infeasibly long. [31] introduced a direct covariance model for Legendre moments, but its final calibration depends on a separate computation for angular bins, which is inconvenient in practice. We extend the methods of [30, 31], developing a more practical and exact estimator for the 2PCF multipole covariance compatible with DESI’s correlation function measurement code, pycorr222https://github.com/cosmodesi/pycorr [33].
@@
The method has been closely integrated with DESI. It has been successfully applied and validated for the early BAO analysis [34, 35]. Following that, it was embraced as part of a coordinated covariance matrix effort for DESI DR1 two-point clustering measurements, along with analytical covariance matrices for power spectra [24] and the general comparison focusing on the consistency of the model fits [36]. We have benefited enormously from synergies with other supporting studies for the galaxies and quasars BAO [7]: optimal reconstruction [37, 38], combined tracers [39], halo occupation distribution systematics [40, 41], fiducial cosmology systematics [42] and theoretical systematics [43].
Fiber assignment incompleteness effects are a big challenge with DESI. For a spectrograph with 5000 robotic fibers, the choice of a target for each is highly complex. Some fibers cannot reach all targets in the field of view, which leads to incomplete coverage, dependent on the priority of different targets and the density of the targets on the sky. To improve the completeness, DESI is scheduled to make 7 dark-time and 4 bright-time passes over each area during its full 5-year program [44]. However, the coverage achieved during the first year of observations (DR1) varies across the footprint [45] between complete and only a single pass in many areas. In multi-pass regions, the assignment also depends on previous DESI exposures in the area. These effects imprint on 2-point clustering [46] and must also affect the higher-point correlations entering the covariance. The exact algorithm can be applied to mocks, but it is not fast enough to obtain as many catalogs () as needed for the estimation or high-quality validation of the full covariance matrix [47]. A faster approximation has been developed [48], enabling a suite of 1000 mocks with fiber assignment modeled. We aim to validate the semi-analytical covariance matrices taking fiber assignment incompleteness into account for the first time.
We organize this paper in the following way. Section 2 reviews the previous results that we build upon: methodology of covariance matrix estimation in Section 2.1, covariance matrix comparison in Section 2.2, and projection to a space of physical model parameters in Section 2.3. In Section 3 we derive the new estimators for covariance matrices of Legendre multipoles of the correlation function, used in the remainder of the paper. Section 4 provides details on the mock catalogs, fiber assignment incompleteness modeling, standard BAO reconstruction and analysis techniques informing our comparison of covariance matrices. Section 5 explains the setup for semi-analytic covariance matrix estimation. In Section 6 we check the consistency of the method’s application to the simulations: intrinsic numerical stability in Section 6.1 and the values of the key parameter of the covariance matrix model in Section 6.2. In Section 7 we compare our semi-analytical covariance matrix estimates with the sample covariance of the mocks in terms of correlation function multipoles in Section 7.1 and the cosmological parameters in Section 7.2, following both BAO and full-shape analyses. Section 8 concludes the main text with a summary and an outlook. Appendix A gives the covariance matrix estimators generalized for multi-tracer analysis. Appendix B explains the procedure to obtain the covariance for the combination of two disjoint regions (volumes).
@@ 2.1 Estimation of covariance matrix for the two-point correlation functions
We start with a summary of the methodology developed in [28, 29, 30, 35]. The RascalC code builds single-parameter333For single tracer; for multiple tracers it is one parameter per tracer. covariance matrix models444This step is the most computationally heavy and is implemented in C++. based on a random catalog and a table of 2-point correlation function values. Then we fit a model to a reference covariance (defined later) to obtain the optimal parameter value for the final prediction. In the fiducial data pipeline, shown schematically in Fig. 1, we measure the correlation function directly from the data, the code produces separate models for full and jackknife covariance matrices, we fit the latter to the data jackknife covariance matrix, and plug the resulting optimal parameter into the full model. Fig. 2 shows an alternative pipeline where we use the best fit of the full covariance model to the mock sample covariance instead. The jackknife methodology has a significant advantage: it only requires mocks for the initial validation, and then allows to generate covariance matrices for different setups without additional simulations. We now provide more details on both pipelines, focusing on the single-tracer case; the generalization to multiple tracers can be found in Appendix A.
@@
To explain the covariance matrix model, we need to start with the definition of the 2-point correlation function (2PCF). The standard Landy-Szalay estimator [49] in radial bin and angular bin 555RascalC assumes uniform binning in , being the cosine of the angle between the line of sight and the pair separation (assuming symmetry with respect to ). The line of sight is the direction to the midpoint of the galaxy pair in an aperiodic survey, and a fixed coordinate direction () in periodic boxes. is
where , are random points and are data points (galaxies).
The random counts are determined by the survey geometry. We assume that the survey design choices are independent of the random realization of the Universe. Consequently, we treat the random counts as fixed.
The numerator, however, is determined by structure formation processes, which are considered stochastic in the cosmological paradigm. Therefore it is crucial for scatter in the correlation function measurements.
The counts can be expanded as
@@
The survey has been divided into cells indexed by and . is the ensemble average number of galaxies in the cell , is the weight for the random point in the cell, is the fractional galaxy overdensity in the cell, is the absolute value of the cosine of the angle between the line of sight and the separation vector , is the length of that vector, and are binning functions (unity if the argument fits into the bin and zero otherwise).
Hereafter we use the following shorthand notation for the covariance matrix:
This covariance matrix involves an ensemble average of 4 overdensities at up to 4 different positions. The reason is that each correlation function estimator (Eqs. 2.1 and 2.2) contains 2 overdensities at 2 different positions, and they are substituted into Eq. 2.3. Some of the 4 positions can be the same, resulting in only 3 or 2 distinct positions666However, all 4 cannot be the same, because the correlation at zero separation is excluded ( in Eq. 2.2). [28].
We then use the following shot-noise approximation to eliminate repeated overdensities at the same position and arrive at the -point correlation functions [28]:
The original motivation [28] refers to the Poisson sampling. We repeat it with slight modifications because the approximation is crucial to the method. Assuming similar weights for galaxies and randoms, the overdensity in cell is
where is the actual number of galaxies in the cell and is the expectation value (ensemble average) of that number777 can vary from cell to cell.. We can choose a sufficiently small cell size so that . Assuming that the galaxies appear in the cell independently, follows the Poisson distribution, and then . Substituting Eq. 2.5 into Eq. 2.4 and taking the ensemble average of both sides gives . This sets the baseline expectation for .
We further explain the overdensity in the right-hand side of Eq. 2.4, because it cancels in the previous argument. With , the average of is dominated by cells with . Whereas almost all the cells have , the corresponding (Eq. 2.5). The fraction of cells with is only , but they have a large . For , the fraction drops significantly to , whereas the does not increase as much. The contributions to the average of (fraction of cells times the value) decrease further for higher values of . Thus substituting the most significant case, , into Eqs. 2.5 and 2.4 gives . This is again true for because . The right-hand side of Eq. 2.4 is dominated by for , therefore this overdensity is necessary.
Several factors can cause the shot-noise rescaling, i.e. effectively shift from 1. First, the observations of galaxies in the same small volume of a real survey are not independent. There are fundamental limitations to their number due to resolution, number of observations, and number and size of optical fibers in a fiber spectroscopic instrument. Furthermore, in DESI each fiber is attached to a robotic positioner confined to a specific area of the focal plane [50]. The selection of a target for the fiber (fiber assignment) therefore must depend on other objects present within this patrol area. As a result, the number of observed galaxies is not guaranteed to follow the Poisson distribution and may not hold. Second, different weights on galaxies and random points can alter the cell overdensity estimate . Such weighting is introduced, in particular, for the mitigation of fiber assignment effects on DESI clustering measurements [46, 45]. The combination of these effects can decrease or increase the shot-noise rescaling.
To recapitulate, the shot-noise approximation (Eq. 2.4) allows us to remove the repeated same-cell overdensities from the covariance matrix estimator. The ensemble averages of a product of overdensities at different positions are the -point correlation function by definition. Therefore the resulting expression for the covariance matrix involves certain sums with the 4-, 3- and 2-point correlation functions. We separate the model covariance matrix into three parts:
@@
These parts, or the -point terms have the following theoretical expressions888In the RascalC code are estimated using Monte Carlo importance sampling of points from random catalogs [30].:
where and are Kronecker deltas; is the 2PCF evaluated999To compute , the RascalC code builds a bicubic interpolator based on an input grid of values. at the separation between points number and .
and are the 3-point and connected 4-point correlation functions. They are evaluated at the separations between and points, respectively. These non-Gaussian higher-point functions are included in the theoretical expression for completeness. However, evaluating them in practice is challenging. Theoretical models may not cover the necessary range of scales and might also require additional assumptions. Direct measurements from data are costly and noisy because the number of bins increases for higher-point correlation functions. Consequently, the 3- and connected 4-point correlation functions are not used in the practical implementation.
Instead of evaluating them, the covariance matrix has been adjusted with the shot-noise rescaling parameter according to Eq. 2.6. The -point terms have been computed solely with the 2-point function (as designated by crossing out and in Eq. 2.7). This 2PCF may, however, include non-linear effects, e.g. due to being estimated from the data or a realistic simulation.
The key shot-noise rescaling parameter value can be chosen for the best match to a reference covariance obtained from a set of mocks. The best match is quantified by minimal Kullback-Leibler divergence101010Which is determined by the covariance matrices for two multivariate normal distributions with the same mean.:
In its computation, the model covariance (with elements given by Eq. 2.6) is inverted because the sample covariance of the simulations is more noisy. With a smooth theoretical model and only one parameter () to adjust, fewer mocks are required than for direct use of their sample covariance matrix [28]. The process is illustrated in Fig. 2. In other words, fitting the RascalC covariance model to mocks is akin to template-based smoothing of the simulation-based sample covariance matrix.
Theoretically, the inversion of the RascalC covariance gives a slightly biased estimate of the precision matrix. However, the Hartlap factor [18] is not applicable since it is not a sample covariance. The relevant correction matrix accounting for importance sampling noise has been worked out [29], but we find it practically insignificant: the eigenvalues deviate from 1 by .
Unfortunately, the mock-fitting approach does not solve all issues with the mocks. The simulations still take extra time to produce and impose further assumptions and approximations. Therefore a different method is desirable to reduce the dependence on simulations.
An alternative reference covariance can be obtained from the data with jackknife resampling. However, the jackknife covariance matrix is not perfectly representative of the true, full-survey covariance for several reasons. First, the resampled pieces of realistic data have different geometry from the full dataset and each other, which affects the covariance in a complicated way. Second, the pieces of the data are correlated [26, 27]. Therefore it is safer to develop a separate theoretical model for jackknife covariance [29], which we briefly explain in the following.
@@
The method uses a slightly non-standard formalism, dubbed unrestricted jackknife [30]. There the jackknife correlation function estimate is not the auto-correlation of the whole survey excluding the jackknife region (as in exclude-one, restricted jackknife), but the cross-correlation function between that region and the whole survey. Equivalently, this means that the additional jackknife weighting factor for the pair of points , , is 1 if both of them belong to the jackknife region , if only one and 0 if neither. The pair counts can be converted from different terms (auto and cross jackknife counts) often saved separately in light of Mohammad-Percival correction [26].
@@
The unrestricted jackknife is convenient since the full pair counts of each type are the sum of all the jackknife pair counts of the same type. Then if one weights the regions by the pair counts,
@@
the weighted mean correlation function is equal to the full-survey one. This simplifies the theoretical modeling of the jackknife covariance.
@@
the corresponding theoretical estimate, , is constructed analogously to Eq. 2.6:
with terms defined as [30]:
where is an additional jackknife weight tensor:
To sum up, in the fiducial (jackknife) pipeline (Fig. 1), we obtain the shot-noise rescaling by fitting the model for its jackknife covariance (Eqs. 2.11 and 2.12) to the data (Eq. 2.10). As in the mock approach, this specifically means minimizing the Kullback-Leibler (KL) divergence between the covariance matrices:
where the model jackknife covariance (Eq. 2.11) is inverted [30], because the data jackknife covariance matrix (Eq. 2.10) is often not invertible. The final covariance is obtained by plugging the resulting shot-noise rescaling values into the full covariance model (Eq. 2.6).
Finally, we need to make adjustments for standard BAO reconstruction because it modifies the correlation function estimator. Standard BAO reconstruction procedures shift the positions of both the data and random points. Only shifted data () is used, whereas the randoms are kept in two variants: original () and shifted (). The correlation function is estimated via the Landy-Szalay estimator (Eq. 2.1) with instead of , but still in the denominator. Following [35], in the computations of covariance matrix terms (Eqs. 2.7 and 2.12) for reconstructed catalogs we use
the shifted randoms in the sums (i.e. used for binning functions and correlation function interpolation are the shifted random positions) because they come from expanding , which now involves and not ;
non-shifted random counts (i.e., still ) for normalization;
two-point correlation functions etc. with a different normalization111111In the code, this is implemented by renormalizing the input grid of correlation values .: having instead of in the denominator, i.e.
For data jackknife covariance (Eq. 2.10), we still use the ordinary normalization of 2PCF (Eq. 2.1 with in the denominator but instead of ).
To summarise, the key assumption is that shot-noise rescaling of purely Gaussian contributions (i.e., ignoring 3-point and connected 4-point functions) can produce a realistic covariance matrix in configuration space. A theoretical motivation is that non-Gaussian contributions primarily affect the squeezed configurations involving small-scale correlations, below the bin width for the 2-point function, therefore not distinguishable from shot noise operating on infinitesimally small scales. The method has been empirically shown to agree well with mock-based covariances [28, 51, 35].